OpenStudy (anonymous):

The radioactive isotope of lead, Pb-209, decays at a rate proportional to the amount present at time t and has a half life of 3.3 hrs. If 1 gram of this isotope is present initially, howlong will it take for 90% of the lead to decay?

OpenStudy (anonymous):

i already know that the proportion is \[\frac{dA}{dt}=kA\]

OpenStudy (anonymous):

dividing by A and multiplying by dt you get \[\frac{dA}{A}=kdt\]

OpenStudy (anonymous):

if you integrate you get \[Ln(A)=kt+c\]

OpenStudy (anonymous):

\[A=ce^{kt}\]

OpenStudy (anonymous):

using the initial condition A(0)=1

OpenStudy (unklerhaukus):

\[c=A_0\]

OpenStudy (anonymous):

yes it is A(0)=c which is 1 \[A=e^{kt}\] correct so far?

OpenStudy (unklerhaukus):

\[90\%=\frac{A}{A_0}=e^{-kt}\]

OpenStudy (anonymous):

the second statement says that after 3.3 years it has a half life is that saying P(3.3)=.5(A(0)

OpenStudy (anonymous):

that's is what i did take the ln of both sides ln(.5)=3.3k

OpenStudy (anonymous):

ln(.5)/3.3=k

OpenStudy (anonymous):

\[A=e^{\frac{ln.5)t}{3.3}}\]

OpenStudy (unklerhaukus):

k or -k, \[\frac{\text dA}{\text dt}=-kA\] for decay

OpenStudy (anonymous):

the book uses kA for hte decay example

OpenStudy (anonymous):

so i assumed that it doens't matter due to it being a constant?

OpenStudy (anonymous):

\[.1=e^{\frac{ln(.5)t}{3.3}}\]

OpenStudy (anonymous):

take ln of both sides? then mutliply by the reciprocal?

OpenStudy (anonymous):

i swear i did the sam ething and got the wrong answer the first time -.- but yes it's correct 10.9 or 11 hrs

OpenStudy (anonymous):

i must have put it wrong in the calculator somehow

OpenStudy (anonymous):

or i tried to simplify the right side way too much

OpenStudy (anonymous):

i think thats what i did i thought the t was added so it was ln(.5)+t and split it

OpenStudy (anonymous):

want to explain kirchoff's law. I was sick the day my teacher expained how it works

OpenStudy (anonymous):

Not an electrical engineer =P lol

OpenStudy (unklerhaukus):

kirchoff's law the sum of current flowing into a junction = the sum of current flowing out of the junction

OpenStudy (anonymous):

yes i mean't the variables R, L, q/c

OpenStudy (unklerhaukus):

what are you asking ?

OpenStudy (anonymous):

\[Ri+\frac{q}{C}=E(t)

OpenStudy (anonymous):

\[Ri+\frac{q}{C}=E(t)\]

OpenStudy (anonymous):

how these formulas are derived. This is probablylike the first time i've ever seen circuits before so i don't know how they relate to eachother

OpenStudy (unklerhaukus):

Driven Harmonic motion is \[m\ddot x+c\dot x+kx=F(t)\] which comes from the Linear Differential Equation \[y''(x)+P(x)y=Q(x)\] In the case of an electric circuit \[L\frac{\text d^2 q}{\text dt^2}+\frac{\text d q}{\text dt}R+\frac qC=V(t)\] if the inductance is zero\[IR+\frac{q}{C}=V(t)\]

OpenStudy (unklerhaukus):

**Second order linear differential equation \[P(x)\frac{\text d^2 y}{\text dx^2}+Q(x)\frac{\text d y}{\text dx}+R(x)y=G(x)\]